If the population size is equal to 5 and the sample size is equal to 3. How many possible sample can be drawn using the formula of without replacement?




1
Expert's answer
2022-05-24T23:05:59-0400

The number of possible samples which can be selected without replacement is

"_NC_n=\cfrac{N! } {n! \cdot(N-n)! };"

"_{5}C_3=\cfrac{5! } {3! \cdot(5-3)! }=\cfrac{4\cdot5}{2}=10."